# Surrogate models for diffusion on graphs via sparse polynomials

### v2

@@ 3.2 Convergence guarantees for least squares and compressed sensing
Theorem 1 will be now leveraged to show algebraic convergence rates for both least squares and compressed sensing (in particular, for weighted SR-LASSO decoders; recall (12)).
Corollary 2 (Convergence of least squares-based surrogates).
Let $0<\epsilon<1$, $\varrho$ be either the uniform or Chebyshev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independent and identically distributed drawn from the probability distribution $\varrho$. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution of equation (4) with the diffusivity coefficient $\boldsymbol{C}(t,\boldsymbol{y})$ satisfying Assumption 2 where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\{\psi_{\boldsymbol{\nu}}\}$ be either the Chebyshev or Legendre basis. Then, there exists a set $S\subset\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f_{v}}=\sum_{\boldsymbol{\nu}\in S}\hat{c}_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$ that solves the least squares problem (9) is unique and satisfies, for every $0<p<1$,
$\|f_{v}-\hat{f_{v}}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq C(\boldsymbol{\rho},p)\cdot(m/\log(m/\epsilon))^{\frac{1}{2}-\frac{1}{p}}+2\cdot\|\boldsymbol{n}\|_{\infty}.$
The proof of Corollary 2, based on [26, Theorem 6.1], can be found in Appendix B.
Theorem 1 also implies uniform and nonuniform guarantees of algebraic convergence (using Theorems 7.12 and 7.13 in [1]) for compressed sensing. In particular, we aim at finding the minimizer of the weighted SR-LASSO decoder defined in (12).
Corollary 3 (Convergence of compressed sensing-based surrogates).
Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution to problem (4) with the diffusivity coefficient $C(t,\boldsymbol{y})$ satisfying Assumption 2, where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})$ be the sequence of coefficients of $f_{v}$ with respect to either the Chebyshev or Legendre basis, Let $m\in\mathbb{N}$ be the number of samples, drawing $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independently from the measure $\varrho$, and $0<\epsilon<1$.
@@
$\begin{split}\|f_{v}-\hat{f_{v}}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq&\;\kappa_{1}\cdot B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})\cdot\big(C_{1}+C_{2}\big)\cdot\tilde{m}^{1/2-1/p}+\kappa_{2}\cdot\|\boldsymbol{n}\|_{2}/\sqrt{m},\\ \|f_{v}-\hat{f_{v}}\|_{L^{\infty}(\mathcal{U})}\leq&\;\kappa_{3}\cdot B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})\cdot\big(C_{1}+C_{2}\big)\cdot\tilde{m}^{1-1/p}+\kappa_{4}\cdot\sqrt{\tilde{m}/m}\cdot\|\boldsymbol{n}\|_{2},\end{split}$
where $\kappa_{i}>0$, $i=1,\dots,4$ are universal constants, $\hat{f}=\sum_{\boldsymbol{\nu}\in\Lambda}\hat{c}_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$, $B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})$ is the costant defined in (22), $C_{1}=C(\boldsymbol{\rho},p,d)$ is as in [1, Lemma 7.19] and $C_{2}=C(\boldsymbol{\rho},p,d)$ is as in [1, Lemma 3.13].
Proof.
The proof follow directly from [1, Theorem 7.12] and Theorem 1. Indeed, Theorem [1, Theorem 7.12] holds when [1, Assumption 2.3] holds. Now, we consider Theorem 1 with the following setting: $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$, $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component), and, consequently, $\mathcal{O}=\mathring{\mathcal{K}}$. Therefore, [1, Assumption 2.3] holds and so does [1, Theorem 7.12]. ∎
@@
Remark 1 (Replacing SR-LASSO with QBCP).
Corollary 3 also holds for weighted QCBP (see (13)). Indeed, [1, Theorem 7.12] is formulated for weighted SR-LASSO, an inspection of its proof reveals that it also holds for weighted QCBP, up to replacing $\|\boldsymbol{n}\|_{2}$ with the QCBP tuning parameter $\eta$ (see also [1, Theorem 7.4], which the proof of [1, Theorem 7.12] crucially relies on).
Remark 2 (From algebraic to exponential rates).
Analogous guarantees of algebraic convergence for uniform recovery, and with exponential decay rates, hold as well. To obtain them it is enough to combine Theorem 1 with [1, Theorem 7.13] and [1, Theorem 7.14] respectively. It is worth observing that the corresponding exponential rates are not dimension independent, though.
Remark 3 (From scalar- to vector-valued solution maps).
@@
which falls under the category of vector-valued compressed sensing (and more generally, compressed sensing for Hilbert-valued functions); see also [1, Section 8.2.5].
@@ 4 Numerical experiments
In this section, we evaluate the efficacy of least squares and compressed sensing for approximating the (real-valued) parametric diffusion map $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{C}(\boldsymbol{y}))$ defined in (19). The primary objective of these experiments is to assess the accuracy of the method across various graph structures, sizes, and diffusion settings. As a performance metric, we use the Root Mean Square Error (RMSE), defined by $\textnormal{RMSE}=\sqrt{\sum_{i=1}^{m_{\text{test}}}\frac{1}{m_{\text{test}}}(f_{v}(\boldsymbol{y}_{i}^{\text{test}})-\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}}))^{2}},$ where $f_{v}(\boldsymbol{y}_{i}^{\text{test}})$ is the actual value and $\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}})$ is the predicted value. In each of the following experiments, the test set will consist of 1000 pairs $(\boldsymbol{y}_{i}^{\text{test}},f(\boldsymbol{y}_{i}^{\text{test}}))$ where the $y_{i}$ are randomly sampled from the uniform distribution on $\mathcal{U}=[-1,1]^{d}$. In this setup we adopt an edge-dependent diffusion, which is a special case of the time-and-edge-dependent diffusion case described by equation (18) in Example 1, where $h_{j}(t)\equiv 1$ for $j\in[d]$. Namely, the matrix $C$ in equation (2) is of the form
@@
In Section 4.1 we first present our experiments on synthetic random graphs, more specifically, generated using the SBM, in order to produce graphs with a community structure. Then, in Section 4.2.1 we replicate some of the same experiments on a real-world graph-structured dataset from Twitter. We implement the proposed method in Python using the NetworkX library for graph generation and manipulation [30], the NumPy library for matrix operations to solve the diffusion equation, and the equadratures [31] and SciPy libraries for the polynomial approximation. The Python code is available at https://github.com/k-yoan/surrogate_graph_diffusion.
4.1 Synthetic graphs
For simplicity, we generate graphs from the SBM such that one graph always has the same number of nodes per community, and the edge probabilities $p_{i,j}$ are always set to $1$ (when nodes $i$ and $j$ belong to the same community) or $0.04$ (when nodes $i$ and $j$ are in different communities). In the subsequent sections, each experiment are repeated 20 times. The geometric mean of RMSE obtained from those rounds is shown in bold colored lines/curves, while the variance is shown with a shaded region around the mean, corresponding to the geometric standard deviation of the RMSE.
@@ 4.1.1 Proof of concept
First, we evaluate the accuracy of least squares and compressed sensing over $m$ sample points, for different dimensions and multi-index sets (total degree and hyperbolic cross, defined in (7) and (8), respectively). We choose the orders $n$ such that the cardinality of the multi-index sets is around the same for both bases ($n=8$ for the total degree set and $n=20$ for the hyperbolic cross). Figure 2 shows that when $d=3$, we attain a high level of accuracy.
Figure 2: Convergence plot of average RMSE vs. number of sample points for $d=3$.
Another thing to note is that when the number of sample points $m$ is less than the cardinality of the basis (represented by the vertical grey dashed line on both plots), QCBP and weighted QCBP yield better results than least squares. This is expected, as least squares only works when the system is overdetermined: as we increase the number of sample points, the average RMSE given by the least squares method decreases and starts performing better than the other two methods.
@@ 4.1.3 Impact of dimension on a fixed size graph
In the experiments, we test for both the total degree set and the hyperbolic cross. For $d=3$, $d=6$ and $d=10$, the cardinalities are 2925, 3003 and 3003, respectively, for the total degree. The results are shown in the left plots of Figures 6, 7 and 8.
@@
Similarly, for the hyperbolic cross, the cardinalities are 3143, 3119 and 3076 (for $d=3$, $d=6$ and $d=10$ respectively). The results are shown in the right plots of Figures 6, 7 and 8. The overall pattern we observe from the three figures is that a lower dimension leads to a better accuracy.
A notable observation is that for higher dimensions, the hyperbolic cross seems to be better suited than the total degree. While the difference is not clearly evident for the least squares case in Figure 6, the QCBP and weighted QCBP cases illustrated in Figures 7 and 8 show that for $d=10$ specifically, the results obtained with the hyperbolic cross are always superior to those with the total degree.

### v3

@@ 3.2 Convergence guarantees for least squares and compressed sensing
Theorem 3.1 will be now leveraged to show algebraic convergence rates for both least squares and compressed sensing (in particular, for weighted SR-LASSO decoders; recall (12)). In particular, the error bounds proved in Corollaries 2 and 3 are “quasi-optimal” in the sense that they follow the same decay rate for the best $s$-term approximation error from Corollary 1 (with $q=2$), i.e., $s^{1/2-1/p}$, after replacing the sparsity $s$ with the number of samples $m$ and up to log factors. Notably, these bounds show that least squares and compressed sensing can converge faster than the Monte Carlo rate $m^{-1/2}$. This is possible thanks to the holomorphic regularity of $f_{v}$. We start by illustrating the least squares case.
Corollary 2 (Convergence of least squares-based surrogates)
Let $0<\epsilon<1$, $\varrho$ be either the uniform or Chebyshev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independent and identically distributed drawn from the probability distribution $\varrho$. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution of equation (4) with the parametric diffusivity matrix $\boldsymbol{C}(t,\boldsymbol{y})$ satisfying Assumption 2 where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\{\psi_{\boldsymbol{\nu}}\}$ be either the Chebyshev or Legendre basis. Then, there exists a set $S\subset\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f_{v}}=\sum_{\boldsymbol{\nu}\in S}\hat{c}_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$ that solves the least squares problem (9) is unique and satisfies, for every $0<p<1$,
$\|f_{v}-\hat{f_{v}}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq B(\boldsymbol{u}_{0},T,\tilde{\boldsymbol{M}},\mathcal{E}_{\boldsymbol{\rho}})\cdot C(\boldsymbol{\rho},p)\cdot(m/\log(m/\epsilon))^{\frac{1}{2}-\frac{1}{p}}+2\cdot\|\boldsymbol{n}\|_{\infty},$
where $B(\boldsymbol{u}_{0},T,\tilde{\boldsymbol{M}},\mathcal{E}_{\boldsymbol{\rho}})$ is as in (22).
The proof of Corollary 2, based on (adcock2022monte, , Theorem 6.1), can be found in Appendix B.
Theorem 3.1 also implies uniform and nonuniform guarantees of algebraic convergence (using Theorems 7.12 and 7.13 in sparsepoly ) for compressed sensing. In particular, we aim at finding the minimizer of the weighted SR-LASSO decoder defined in (12).
Corollary 3 (Convergence of compressed sensing-based surrogates)
Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution to problem (4) with the parametric diffusivity matrix $\boldsymbol{C}(t,\boldsymbol{y})$ satisfying Assumption 2, where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})$ be the sequence of coefficients of $f_{v}$ with respect to either the Chebyshev or Legendre basis, Let $m\in\mathbb{N}$ be the number of samples, drawing $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independently from the measure $\varrho$, and $0<\epsilon<1$.
@@
$\begin{split}\|f_{v}-\hat{f_{v}}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq&\;\kappa_{1}\cdot B(\boldsymbol{u}_{0},T,\tilde{\boldsymbol{M}},\mathcal{E}_{\boldsymbol{\rho}})\cdot\big(C_{1}+C_{2}\big)\cdot\tilde{m}^{1/2-1/p}+\kappa_{2}\cdot\|\boldsymbol{n}\|_{2}/\sqrt{m},\\ \|f_{v}-\hat{f_{v}}\|_{L^{\infty}(\mathcal{U})}\leq&\;\kappa_{3}\cdot B(\boldsymbol{u}_{0},T,\tilde{\boldsymbol{M}},\mathcal{E}_{\boldsymbol{\rho}})\cdot\big(C_{1}+C_{2}\big)\cdot\tilde{m}^{1-1/p}+\kappa_{4}\cdot\sqrt{\tilde{m}/m}\cdot\|\boldsymbol{n}\|_{2},\end{split}$
where $\kappa_{i}>0$, $i=1,\dots,4$ are universal constants, $\hat{f}=\sum_{\boldsymbol{\nu}\in\Lambda}\hat{c}_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$, $B(\boldsymbol{u}_{0},T,\tilde{\boldsymbol{M}},\mathcal{E}_{\boldsymbol{\rho}})$ is the constant defined in (22), $C_{1}=C(\boldsymbol{\rho},p,d)$ is as in (sparsepoly, , Lemma 7.19) and $C_{2}=C(\boldsymbol{\rho},p,d)$ is as in (sparsepoly, , Lemma 3.13).
Proof
The proof follow directly from (sparsepoly, , Theorem 7.12) and Theorem 3.1. Indeed, Theorem (sparsepoly, , Theorem 7.12) holds when (sparsepoly, , Assumption 2.3) holds. Now, we consider Theorem 3.1 with the following setting: $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$, $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component), and, consequently, $\mathcal{O}=\mathring{\mathcal{K}}$. Therefore, (sparsepoly, , Assumption 2.3) holds and so does (sparsepoly, , Theorem 7.12).
@@
Remark 3(Replacing SR-LASSO with QBCP)
Corollary 3 also holds for weighted QCBP (see (13)). Indeed, (sparsepoly, , Theorem 7.12) is formulated for weighted SR-LASSO, an inspection of its proof reveals that it also holds for weighted QCBP, up to replacing $\|\boldsymbol{n}\|_{2}$ with the QCBP tuning parameter $\eta$ (see also (sparsepoly, , Theorem 7.4), which the proof of (sparsepoly, , Theorem 7.12) crucially relies on).
Remark 4(From algebraic to exponential rates)
Analogous guarantees of algebraic convergence for uniform recovery, and with exponential decay rates, hold as well. To obtain them it is enough to combine Theorem 3.1 with (sparsepoly, , Theorem 7.13) and (sparsepoly, , Theorem 7.14) respectively. It is worth observing that the corresponding exponential rates are not dimension independent, though.
Remark 5(From scalar- to vector-valued solution maps)
@@
which falls under the category of vector-valued compressed sensing (and more generally, compressed sensing for Hilbert-valued functions); see also (sparsepoly, , Section 8.2.5).
@@ 4 Numerical experiments
In this section, we evaluate the efficacy of least squares and compressed sensing for approximating the (real-valued) parametric diffusion map $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{C}(\boldsymbol{y}))$ defined in (19). The primary objective of these experiments is to assess the accuracy of the method across various graph structures, sizes, and diffusion settings. As a performance metric, we use the Root Mean Square Error (RMSE), defined by $\textnormal{RMSE}=\sqrt{\sum_{i=1}^{m_{\text{test}}}\frac{1}{m_{\text{test}}}(f_{v}(\boldsymbol{y}_{i}^{\text{test}})-\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}}))^{2}},$ where $f_{v}(\boldsymbol{y}_{i}^{\text{test}})$ is the actual value and $\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}})$ is the predicted value. In each of the following experiments, the test set will consist of 1000 pairs $(\boldsymbol{y}_{i}^{\text{test}},f(\boldsymbol{y}_{i}^{\text{test}}))$ where the $\boldsymbol{y}_{i}$’s are randomly sampled from the uniform distribution on $\mathcal{U}=[-1,1]^{d}$.
In Section 4.1 we first present our experiments on synthetic random graphs, more specifically, generated using the SBM, in order to produce graphs with a community structure. We then increase the complexity of the case study by focusing on a time-and-edge-dependent diffusion problem in Section 4.2. In Section 4.3 we replicate some of the same experiments of Section 4.1 on real-world graph-structured datasets from Twitter and Facebook. Finally, in Section 4.4 we carry out a computational cost analysis of the proposed surrogates compared to full-order simulations. We implement the proposed method in Python using the NetworkX library for graph generation and manipulation hagberg2008exploring , the NumPy library for matrix operations to solve the diffusion equation, and the equadratures equad and SciPy libraries for the polynomial approximation. The Python code is available at https://github.com/k-yoan/surrogate_graph_diffusion.
We conclude with a general remark regarding numerical comparisons between least squares and compressed sensing. While this section includes several experiments where approximations obtained using these two techniques are compared directly, it is important to recall that these two methods are designed to be used in different sampling regimes. Indeed, on the one hand least squares is only able to provide meaningful approximations in the oversampled regime. On the other hand, compressed sensing works in the undersampled regime, but it becomes suboptimal in the oversamples regime due to the presence of bias induced by sparse regularization. Therefore, the numerical experiments where both techniques are applied in different regimes should not be thought of as an algorithmic comparison aimed at establishing a “winner”, but rather as qualitative illustrations showcasing the behavior of these methods in different scenarios.
4.1 Synthetic graphs (edge-dependent case)
In this section we consider an edge-dependent diffusion, which is a special case of the time-and-edge-dependent diffusion case described by equation (18) in Example 1, where $h_{j}(t)\equiv 1$ for $j\in[d]$. Namely, the matrix $\boldsymbol{C}$ in equation (2) is of the form
@@ 4.1 Synthetic graphs (edge-dependent case)
For simplicity, we generate graphs from the SBM such that one graph always has the same number of nodes per community, and the edge probabilities $p_{i,j}$ are always set to $1$ (when nodes $i$ and $j$ belong to the same community) or $0.04$ (when nodes $i$ and $j$ are in different communities). In the subsequent sections, each experiment is repeated 20 times. The geometric mean of RMSE obtained from those rounds is shown in bold colored lines/curves, while the variance is shown with a shaded region around the mean, corresponding to the geometric standard deviation of the RMSE.
For the moment, we will focus on time-independent problems. This choice is made mainly for computational simplicity. In fact, our numerical experiments involve statistical simulations based on Monte Carlo-type experiments, where independent experiments are run multiple times. In the time independent case, computing samples is very fast since we rely on an explicit forumla that depends on the eigendecomposition of the graph Laplacian $\boldsymbol{L}$. On the other hand, in the case of time-dependent problems samples are computed via a much slower numerical integration. Specifically, although Sections 4.1.14.1.4 focus on time-independent problems, Sections 4.24.4 will include time-dependent problems.
@@ 4.1.1 Proof of concept
In order to numerically validate our theoretical findings, we start by studying the decay and summability properties of the polynomial coefficients of the parametric diffusion map $f_{v}(\boldsymbol{y})$. In particular, we compute its polynomial coefficients in a total degree set of order $n=10$ using least squares with a large number of samples (namely, $m=10000$) to guarantee very high accuracy. Then, we approximate $\sigma_{s}(\boldsymbol{c})_{2}$ of Definition 1 with the best $s$-term approximation error of the computed coefficient vector. The results are shown in Figure 2. We can see that the sorted absolute coefficients (Figure 2, left) and the best $s$-term approximation error exhibit a super-algebraic decay. This is due to the holomorphic regularity of the map, proved in Theorem 3.1, and is a direct consequence of (sparsepoly, , Theorem 3.2) and Corollary 1, respectively.
Figure 2: Absolute polynomial coefficients of the parametric diffusion map $f_{y}(\boldsymbol{y})$ (left: unsorted, center: sorted) and best $s$-term approximation error (right).
Now, we evaluate the accuracy of least squares and compressed sensing over $m$ sample points, for different dimensions and multi-index sets (total degree and hyperbolic cross, defined in (7) and (8), respectively). We choose the orders $n$ such that the cardinality of the multi-index sets is around the same for both bases ($n=8$ for the total degree set and $n=20$ for the hyperbolic cross). Figure 3 shows that when $d=3$, we attain a high level of accuracy.
Figure 3: Convergence plot of average RMSE vs. number of sample points for $d=3$. The green and red dashed lines highlight the order of convergence of the methods. The vertical black dashed line represents the transition from the underdetermined to the overdetermined regime, i.e., $m=N$.
Another thing to note is that when the number of sample points $m$ is less than the cardinality of the basis (represented by the vertical grey dashed line on both plots), QCBP and weighted QCBP yield better results than least squares. This is expected, as least squares only works when the system is overdetermined: as we increase the number of sample points, the average RMSE given by the least squares method decreases and starts performing better than the other two methods. In addition, we compare the observed convergence plots with curves of the form $(\log(m)/m)^{1/2-1/p}$ for different values of $p$ (see Corollaries 2 and 3). In all cases we find that the observed converges qualitatively exhibits this behavior for some value of $0<p<1$. This illustrates that it is possible to numerically observe convergence rates of the form predicted by Corollaries 2 and 3 (modulo a simpler log factor in the latter case), which are faster than the Monte Carlo rate $m^{-1/2}$.
@@ 4.1.3 Impact of dimension on a fixed size graph
In the experiments, we test for both the total degree set and the hyperbolic cross. For $d=3$, $d=6$ and $d=10$, the cardinalities are 2925, 3003 and 3003, respectively, for the total degree. The results are shown in the left plots of Figures 7, 8 and 9.
@@
Similarly, for the hyperbolic cross, the cardinalities are 3143, 3119 and 3076 (for $d=3$, $d=6$ and $d=10$ respectively). The results are shown in the right plots of Figures 7, 8 and 9. The overall pattern we observe from the three figures is that a lower dimension leads to a better accuracy.
A notable observation is that for higher dimensions, the hyperbolic cross seems to be better suited than the total degree. While the difference is not clearly evident for the least squares case in Figure 7, the QCBP and weighted QCBP cases illustrated in Figures 8 and 9 show that for $d=10$ specifically, the results obtained with the hyperbolic cross are always superior to those with the total degree.